# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ 3.1. Convergence of One-point Distributions
$\hat{\pi}_{s,t}^{(N),d,d}\left(\hat{y}_{j}^{(N),u}(s),\hat{y}_{k}^{(N),u}(t)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{j}^{(N),u}(s)\right)\frac{\pi_{q^{t}}^{(N),d}\left(y_{k}^{(N),u}(t)\right)}{\pi_{q^{s}}^{(N),d}\left(y_{j}^{(N),u}(s)\right)}$
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t}{2}-s+u+k},q^{\frac{t}{2}-u-k}\right)$
which also has no atoms from $u-\frac{s}{2}<0$. ∎
@@
$\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),1-\frac{r}{2N}\right)$ $=\frac{1}{4\pi N\sqrt{\frac{r}{N}}\sqrt{1-\frac{r}{4N}}}\times$
$\frac{(q,-q^{v+w+s},q^{t+j+2v},q^{-j},-q^{t+w+j+v},-q^{w-j-v},q^{t-s})_{\infty}|(q^{i\sqrt{N}\theta_{r}})|^{2}}{(-q^{v+w+t})_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}+(j+v+\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-(j+v+\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}}\right)_{\infty}\bigg{|}^{2}}$
@@
$\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),u}(t),1-\frac{r}{2N}\right)$ $=\frac{1}{4\pi N\sqrt{\frac{r}{N}}\sqrt{1-\frac{r}{4N}}}\times$
$\frac{(q,-q^{v+w+s},q^{j+v+u},q^{v-j-u+t},-q^{w+j+u},-q^{w-j-u+t},q^{t-s})_{\infty}|(q^{i\sqrt{N}\theta_{r}})|^{2}}{(-q^{v+w+t})_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}+(j+u-\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-(j+u-\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}}\right)_{\infty}\bigg{|}^{2}}$
@@
$\sqrt{1-\frac{r}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),u}(t),1-\frac{r}{2N}\right)=$
$\frac{\big{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(j+u-\frac{s}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(t-\frac{s}{2}-j-u+i\frac{\sqrt{r}}{2}\right)\big{|}^{2}}{8\pi\sqrt{r}\Gamma\left(j+v+u\right)\Gamma\left(v-j-u+t\right)\Gamma\left(t-s\right)|\Gamma(i\sqrt{r})|^{2}}\cdot e^{\textsc{Error}_{t,s}^{(N)}(y^{(N),u}_{j}(t))}$
with the error term bounded above for all $r\in\hat{S}^{(N),c}=[0,4N]$ as
@@
4. Proof of Proposition 3.1
In this section, we put the results from Section 3 together to prove 1.3 $(2)$. We begin by defining functions
$\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})$ $:=\mathbf{1}_{\textbf{r}\in\mathbb{R}^{d+1}_{\leq 4N}}\cdot 2^{-N}\prod_{k=1}^{d+1}\left(\cosh{\left(\frac{s_{k}}{\sqrt{N}}\right)}+1-\frac{r_{k}}{2N}\right)^{N(x_{k}^{(N)}-x_{k-1}^{(N)})}.$
$\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})$ $:=\exp{\left(\frac{1}{4}\sum_{k=1}^{d}(s_{k}^{2}-r_{k})(x_{k}-x_{k-1})\right)}$
We make these definitions in order to concisely rewrite the formula for the Laplace transform of the diffusively scaled height function (2.1) in terms of $\mathcal{G}^{(N)}(\mathbf{r};\mathbf{c};\mathbf{x})$ and $\mathcal{G}(\mathbf{r};\mathbf{c};\mathbf{x})$,
$\varphi^{(N)}_{u,v,w,r}(\textbf{c},\textbf{x})$ $\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh{\left(s_{k}^{(N)}\right)}+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]},$
$=\frac{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{s_{1}}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{s_{d+1}});\textbf{c};\textbf{x}\right)\right]}{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{0}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{0});\mathbf{0};\textbf{x}\right)\right]}=\frac{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{s_{1}}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{s_{d+1}});c;X\right)\right]}{2^{-N}\mathbb{E}\left[\left(1-\frac{\hat{\mathbb{Y}}_{0}^{(N)}}{4N}\right)^{N}\right]}.$
$\varphi_{u,v}(\textbf{c};\textbf{x})$ $:=\frac{\mathbb{E}\left[\exp{\left(\frac{1}{4}\sum_{k=1}^{d}\left(s_{k}^{2}-\mathbb{T}_{s_{k}}\right)\left(x_{k}-x_{k-1}\right)\right)}\right]}{\mathbb{E}\left[\left(-\frac{1}{4}\mathbb{T}_{0}\right)^{N}\right]}$
$=\frac{\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{s_{1}},...,\mathbb{T}_{s_{d}});\textbf{c};\textbf{x}\right)\right]}{\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{0},...,\mathbb{T}_{0});\textbf{c};\textbf{x}\right)\right]}.$
@@
$\widetilde{\varphi}_{u,v,w,r}^{(N)}(\textbf{c};\textbf{x})$ $:=\int\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\prod_{k=1}^{d}\hat{\pi}^{(N)}_{s_{i+1},s_{i}}(r_{i+1},dr_{i})\cdot N^{u+v}\hat{\pi}^{(N)}_{s_{d+1}}(dr_{d+1})$
$=N^{u+v}\mathbb{E}\left[\mathcal{G}^{(N)}\left((\hat{\mathbb{Y}}_{s_{1}}^{(N)},...,\hat{\mathbb{Y}}^{(N)}_{s_{d+1}});\textbf{c};\textbf{x}\right)\right]$
$\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x})$ $:=\int\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})\prod_{k=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}(r_{i+1},dr_{i})\cdot\mathfrak{p}_{s_{d+1}}(dr_{d+1})$
$=\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{s_{1}},...,\mathbb{T}_{s_{d}});\textbf{c};\textbf{x}\right)\right]$
@@
$\varphi^{(N)}_{u,v,w,r}(\textbf{c};\textbf{x})=\frac{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\textbf{c};\textbf{x})}{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\mathbf{0};\textbf{x})}.$
The proof of 1.3 $(2)$ will proceed by demonstrating that (Lemma 4.3)
$\lim_{N\to\infty}\widetilde{\varphi}_{u,v,w,r}^{(N)}(\textbf{c};\textbf{x})=C_{u,v,w,r}\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x}).$
We cite the following lemma from Corwin and Knizel [CK24].
@@
For all $N\in\mathbb{N}$, the functions $\widetilde{\varphi}_{u,v}(\textbf{c},\textbf{x})$, $\textbf{c}\in(0,C_{u,r})^{d}$, and $\widetilde{\varphi}_{u,v}(\textbf{0},\textbf{x})$ take values in $(0,\infty)$.
The proof is identical to that in [CK24], so we only paraphrase the argument here. The first step is to apply the bound $\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})\leq Ce^{-(1-x_{d})r_{d+1}/4}$ and then repeatedly apply identities in Lemma 2.4 to integrate over the variables $r_{i}$ to obtain
$\widetilde{\varphi}_{v,w,u,r}(\textbf{c};\textbf{x})\leq C\int_{\mathbb{R}}e^{-(1-x_{d})r}\mathfrak{p}_{0}(dr).$
Then, it is necessary to apply further bounds to the asymptotic behavior of $\mathfrak{p}_{0}^{c}$ to show that this integral is finite. Showing that $\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x})$ is positive involves noting that $\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})\geq\mathcal{G}(\mathbf{2};\textbf{c};\textbf{x})$ for all $\mathbf{r}\in[1,2]^{d+1}$ and then using this bound to obtain the lower bound
$\widetilde{\varphi}_{u,v}(\mathbf{c};\mathbf{x})\geq\mathcal{G}(\mathbf{2};\mathbf{c};\mathbf{x})\int_{[1,2]^{d+1}}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{c,c}(r_{i+1},r_{i})dr_{i}\cdot\mathfrak{p}_{s_{d+1}}^{c}(r_{d+1})dr_{d+1}$
@@
For every compact interval $I\subset\mathbb{R}$, there exists a constant $C>0$ such that for all $\textbf{c}\in\mathbb{R}^{d+1}$,
$\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\leq C\prod_{k=1}^{d+1}e^{-(x_{k}-x_{k-1})\frac{r_{k}}{4}+\frac{|r_{k}|}{2N}}$
where x and c are defined in (2.2). Furthermore, for all $\textbf{r}^{(N)}\in\mathbb{R}^{d+1}$ with $\lim_{N\to\infty}\textbf{r}^{(N)}=\textbf{r},$
$\lim_{N\to\infty}\mathcal{G}^{(N)}(\textbf{r}^{(N)};\textbf{c};\textbf{x})=\mathcal{G}(\textbf{r};\textbf{c};\textbf{x})$
@@
Finally, we will show the convergence of the sequence $\{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\textbf{c};\textbf{x})\}_{N\in\mathbb{N}}$.
@@
For $(q,A,B,C,D)$ and $h_{u,v,w,r}^{(N)}(x)$ satisfying the scaling conditions defined in 1.1, and for all $\textbf{c}\in(0,C_{u,r})^{d}$ and $\textbf{c}=\textbf{0}$,
$\lim_{N\to\infty}\widetilde{\varphi}_{u,v,w,r}^{(N)}(\textbf{c};\textbf{x})=C_{u,v,w,r}\widetilde{\varphi}_{u,v}(\textbf{c};\textbf{x}).$
The argument is unchanged from that given in Corwin and Knizel’s paper [CK24], albeit using the lemmas we established in Section 3 for general $w$ and $r$ and the scaling conditions in 1.1. For this reason, we only provide a summary of the argument here.
@@
(4.1) $\bigg{|}\mathcal{G}^{(N)}(\textbf{r};\textbf{c};\textbf{x})\prod_{i=1}^{d}\hat{\pi}_{s_{i+1},s_{i}}^{(N),a_{i+1},a_{i}}(r_{i+1},r_{i})\cdot\hat{\pi}_{s_{d+1}}^{(N),a_{d+1}}(r_{d+1})\bigg{|},$
$\leq Ce^{-c(r_{1}+...+r_{d+1}}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{a_{i+1},a_{i}}\cdot\mathfrak{p}_{s_{d+1}}^{a_{d+1}}(r_{d+1}).$
All that remains is to show that, for all $I\subset[[1,d+1]]$,
@@
which is accomplished by casework on the structure of the set $I$.
5. Proof of Theorem 1.3
In this section, we discuss the tightness of the laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$, and then give a full proof of 1.3. A standard result in the study of stochastic processes (see, for instance, Karatzas and Shreve [KS91]) tells us that if, for all $0\leq x,y\leq 1$ there exist $\alpha,\beta,\nu,C\in\mathbb{R}_{>0}$ such that
(5.1) $\sup_{N\in\mathbb{N}}\mathbb{E}\left[|h_{u,v,w,r}^{(N)}(x)-h_{u,v,w,r}^{(N)}(y)|^{\alpha}\right]\leq C|x-y|^{1+\beta},$
and such that the process, evaluated at 00, has bounded supremum (which is immediate for our process since $h^{(N)}_{u,v,w,r}(0)-h^{(N)}_{u,v,w,r}(0)=0$), then the collection of laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ is tight on $(C[0,1],\mathcal{B}(C[0,1]))$. Since $h_{u,v,w,r}^{(N)}(0)-h_{u,v,w,r}^{(N)}(0)=0$ for all $N\in\mathbb{N}$, the proof of tightness reduces to demonstrating the second bound in (5.1).

### v2

@@ 3.1. Convergence of Marginal Distributions
$\widehat{\pi}_{s,t}^{(N),d,d}\left(\widehat{y}_{j}^{(N),u}(s),\widehat{y}_{k}^{(N),u}(t)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{j}^{(N),u}(s)\right)\frac{\pi_{q^{t}}^{(N),d}\left(y_{k}^{(N),u}(t)\right)}{\pi_{q^{s}}^{(N),d}\left(y_{j}^{(N),u}(s)\right)}$
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+u+k-\frac{t}{2}},q^{\frac{t0s}{2}-(u+k-\frac{t}{2})}\right),$
which no atoms associated to the first and second arguments, due to the constraint $u+v>0$ and $u-s/2<0$ as well as the restriction of the domain to avoid atoms arising from $w+s/2<0.$ Only the third argument can attain atoms, which arise whenever $-s/2+u+k<0$, in which case the term will contribute an atom $y^{(N),u}_{k+i}(s)$ for all $i\in[[0,\lfloor-u-k+\frac{s}{2}\rfloor]]$. The masses at these atoms are given by
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{k}^{(N),u}(s)\right)$ $=\frac{(q^{-2u-2k+s},-q^{v+1+s},q^{v-u-k+t},-q^{1-u-k+t})_{\infty}}{(q^{v-u-k+s},-q^{1-u-k+s},q^{-2u-2k+t},-q^{v+t+1})_{\infty}},$
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{k+i}^{(N),u}(s)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{k}^{(N),u}(s)\right)\cdot M_{i}$
$M_{i}$ $:=\frac{(q^{2u+2k-s},q^{v+u+k},-q^{u+1+k},q^{t-s})_{i}(1-q^{2u+2k+2i-s})}{(q,q^{u-v+1+k-s},-q^{u+k-s},q^{2u+2k-t+1})_{i}(1-q^{2u+2k-s})}\cdot(-q^{-v-2t+s})^{i}$
Under the same analysis as in preceding discussions, we see that
$\lim_{N\to\infty}\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),u}(t),y_{k}^{(N),u}(s)\right)$ $=\frac{\Gamma(v-u-k+s)\Gamma(-2u-2k+t)}{\Gamma(-2u-2k+s)\Gamma(v-u-k+t)}$
$\lim_{N\to\infty}M_{i}$ $=\frac{[2u+2k-s,v+u+k,t-s]_{i}(2u+2k+2i-s)}{[1,u-v+1+k-s,2u+2k-t+1]_{i}(2u+2k-s)}(-1)^{i}.$
Putting this together, and replicating the asymptotic bounds from previous computations finishes the proof of Lemma 3.3 (2). ∎
@@
$\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),1-\frac{r}{2N}\right)$ $=\frac{1}{4\pi N\sqrt{\frac{r}{N}}\sqrt{1-\frac{r}{4N}}}$
×$\times\frac{(q,-q^{v+w+s},q^{t+j+2v},q^{-j},-q^{t+w+j+v},-q^{w-j-v},q^{t-s})_{\infty}|(q^{i\sqrt{N}\theta_{r}})|^{2}}{(-q^{v+w+t})_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}+(j+v+\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-(j+v+\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}}\right)_{\infty}\bigg{|}^{2}}.$
@@
$\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),u}(t),1-\frac{r}{2N}\right)$ $=\frac{1}{4\pi N\sqrt{\frac{r}{N}}\sqrt{1-\frac{r}{4N}}}$
×$\times\frac{(q,-q^{v+w+s},q^{j+v+u},q^{v-j-u+t},-q^{w+j+u},-q^{w-j-u+t},q^{t-s})_{\infty}|(q^{i\sqrt{N}\theta_{r}})|^{2}}{(-q^{v+w+t})_{\infty}\bigg{|}\left(q^{v+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}+(j+u-\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-(j+u-\frac{t}{2})+i\frac{\sqrt{N}\theta_{r}}{2}}\right)_{\infty}\bigg{|}^{2}}.$
@@
$\sqrt{1-\frac{r}{4N}}\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),u}(t),1-\frac{r}{2N}\right)$
$=\frac{\big{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(j+u-\frac{s}{2}+i\frac{\sqrt{r}}{2}\right)\Gamma\left(t-\frac{s}{2}-j-u+i\frac{\sqrt{r}}{2}\right)\big{|}^{2}}{8\pi\sqrt{r}\Gamma\left(j+v+u\right)\Gamma\left(v-j-u+t\right)\Gamma\left(t-s\right)|\Gamma(i\sqrt{r})|^{2}}\cdot e^{\textsc{Error}_{t,s}^{(N)}(y^{(N),u}_{j}(t))},$
with the error term bounded above for all $r\in\widehat{S}^{(N),c}=[0,4N]$ as
@@
4. Convergence of the Laplace Transform
In this section, we put the results from Section 3 together to prove 3.1. We begin by defining functions
$\mathcal{G}^{(N)}(\vec{r};\vec{c};\vec{x})$ $:=\mathbf{1}_{\vec{r}\in\mathbb{R}^{d+1}_{\leq 4N}}\cdot 2^{-N}\prod_{k=1}^{d+1}\left(\cosh{\left(\frac{s_{k}}{\sqrt{N}}\right)}+1-\frac{r_{k}}{2N}\right)^{N(x_{k}^{(N)}-x_{k-1}^{(N)})},$
$\mathcal{G}(\vec{r};\vec{c};\vec{x})$ $:=\exp{\left(\frac{1}{4}\sum_{k=1}^{d}(s_{k}^{2}-r_{k})(x_{k}-x_{k-1})\right)}.$
We make these definitions in order to concisely rewrite the formula for the Laplace transform of the diffusively scaled height function (2.1) in terms of $\mathcal{G}^{(N)}(\vec{r};\vec{c};\vec{x})$ and $\mathcal{G}(\vec{r};\vec{c};\vec{x})$,
$\varphi^{(N)}_{u,v,w,r}(\vec{c},\vec{x})$ $:=\frac{\mathbb{E}\left[\prod_{k=1}^{d+1}\left(\cosh{\left(s_{k}^{(N)}\right)}+\mathbb{Y}_{e^{-2s_{k}^{(N)}}}\right)^{n_{k}-n_{k-1}}\right]}{\mathbb{E}\left[(1+\mathbb{Y}_{1})^{N}\right]}$
$=\frac{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\widehat{\mathbb{Y}}_{s_{1}}^{(N)},...,\widehat{\mathbb{Y}}^{(N)}_{s_{d+1}});\vec{c};\vec{x}\right)\right]}{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\widehat{\mathbb{Y}}_{0}^{(N)},...,\widehat{\mathbb{Y}}^{(N)}_{0});\vec{0};\vec{x}\right)\right]}=\frac{\mathbb{E}\left[\mathcal{G}^{(N)}\left((\widehat{\mathbb{Y}}_{s_{1}}^{(N)},...,\widehat{\mathbb{Y}}^{(N)}_{s_{d+1}});c;X\right)\right]}{2^{-N}\mathbb{E}\left[\left(1-\frac{\widehat{\mathbb{Y}}_{0}^{(N)}}{4N}\right)^{N}\right]},$
$\varphi_{u,v}(\vec{c};\vec{x})$ $:=\frac{\mathbb{E}\left[\exp{\left(\frac{1}{4}\sum_{k=1}^{d}\left(s_{k}^{2}-\mathbb{T}_{s_{k}}\right)\left(x_{k}-x_{k-1}\right)\right)}\right]}{\mathbb{E}\left[\left(-\frac{1}{4}\mathbb{T}_{0}\right)^{N}\right]}$
$=\frac{\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{s_{1}},...,\mathbb{T}_{s_{d}});\vec{c};\vec{x}\right)\right]}{\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{0},...,\mathbb{T}_{0});\vec{c};\vec{x}\right)\right]}.$
@@ 4. Convergence of the Laplace Transform
$\widetilde{\varphi}_{u,v,w,r}^{(N)}(\vec{c};\vec{x})$ $:=\int\mathcal{G}^{(N)}(\vec{r};\vec{c};\vec{x})\prod_{k=1}^{d}\widehat{\pi}^{(N)}_{s_{i+1},s_{i}}(r_{i+1},dr_{i})\cdot N^{u+v}\widehat{\pi}^{(N)}_{s_{d+1}}(dr_{d+1})$
$=N^{u+v}\mathbb{E}\left[\mathcal{G}^{(N)}\left((\widehat{\mathbb{Y}}_{s_{1}}^{(N)},...,\widehat{\mathbb{Y}}^{(N)}_{s_{d+1}});\vec{c};\vec{x}\right)\right]$
$\widetilde{\varphi}_{u,v}(\vec{c};\vec{x})$ $:=\int\mathcal{G}(\vec{r};\vec{c};\vec{x})\prod_{k=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}(r_{i+1},dr_{i})\cdot\mathfrak{p}_{s_{d+1}}(dr_{d+1})$
$=\mathbb{E}\left[\mathcal{G}\left((\mathbb{T}_{s_{1}},...,\mathbb{T}_{s_{d}});\vec{c};\vec{x}\right)\right].$
@@
$\varphi^{(N)}_{u,v,w,r}(\vec{c};\vec{x})=\frac{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\vec{c};\vec{x})}{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\vec{0};\vec{x})}.$
The proof of 1.1 will proceed by demonstrating that (Lemma 4.3)
$\lim_{N\to\infty}\widetilde{\varphi}_{u,v,w,r}^{(N)}(\vec{c};\vec{x})=C_{u,v,w,r}\widetilde{\varphi}_{u,v}(\vec{c};\vec{x}).$
We cite the following lemma from [CK24].
@@
For all $N\in\mathbb{N}$, the functions $\widetilde{\varphi}_{u,v}(\vec{c},\vec{x})$, $\vec{c}\in(0,C_{u,r})^{d}$, and $\widetilde{\varphi}_{u,v}(\vec{0},\vec{x})$ take values in $(0,\infty)$.
The proof is identical to that in [CK24], so we only paraphrase the argument here. The first step is to apply the bound $\mathcal{G}(\vec{r};\vec{c};\vec{x})\leq Ce^{-(1-x_{d})r_{d+1}/4}$ and then repeatedly integrate over the variables $r_{i}$ to obtain
$\widetilde{\varphi}_{v,w,u,r}(\vec{c};\vec{x})\leq C\int_{\mathbb{R}}e^{-(1-x_{d})r}\mathfrak{p}_{0}(dr).$
Then, it is necessary to apply further bounds to the asymptotic behavior of $\mathfrak{p}_{0}^{c}$ to show that this integral is finite. Showing that $\widetilde{\varphi}_{u,v}(\vec{c};\vec{x})$ is positive involves noting that $\mathcal{G}(\vec{r};\vec{c};\vec{x})\geq\mathcal{G}(\vec{2};\vec{c};\vec{x})$ for all $\vec{r}\in[1,2]^{d+1}$ and then using this bound to obtain the lower bound
$\widetilde{\varphi}_{u,v}(\vec{c};\vec{x})\geq\mathcal{G}(\vec{2};\vec{c};\vec{x})\int_{[1,2]^{d+1}}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{c,c}(r_{i+1},r_{i})dr_{i}\cdot\mathfrak{p}_{s_{d+1}}^{c}(r_{d+1})dr_{d+1}.$
@@
For every compact interval $I\subset\mathbb{R}$, there exists a constant $C>0$ such that for all $\vec{c}\in\mathbb{R}^{d+1}$,
$\mathcal{G}^{(N)}(\vec{r};\vec{c};\vec{x})\leq C\prod_{k=1}^{d+1}e^{-(x_{k}-x_{k-1})\frac{r_{k}}{4}+\frac{|r_{k}|}{2N}},$
where $\vec{x}$ and $\vec{c}$ are defined in (2.2). Furthermore, for all $\vec{r}^{(N)}\in\mathbb{R}^{d+1}$ with $\lim_{N\to\infty}\vec{r}^{(N)}=\vec{r},$
$\lim_{N\to\infty}\mathcal{G}^{(N)}(\vec{r}^{(N)};\vec{c};\vec{x})=\mathcal{G}(\vec{r};\vec{c};\vec{x}).$
@@
Finally, we will show the convergence of the sequence $\{\widetilde{\varphi}^{(N)}_{u,v,w,r}(\vec{c};\vec{x})\}_{N\in\mathbb{N}}$.
@@
For $(q,A,B,C,D)$ and $h_{u,v,w,r}^{(N)}(x)$ satisfying the scaling conditions defined in 1.1, and for all $\vec{c}\in(0,C_{u,r})^{d}$ and $\vec{c}=\vec{0}$,
$\lim_{N\to\infty}\widetilde{\varphi}_{u,v,w,r}^{(N)}(\vec{c};\vec{x})=C_{u,v,w,r}\widetilde{\varphi}_{u,v}(\vec{c};\vec{x}).$
The argument is unchanged from that given in [CK24], albeit using the lemmas we established in Section 3 for general $w$ and $r$ and the scaling conditions in 1.1. For this reason, we only provide a summary of the argument here.
@@
$\bigg{|}\mathcal{G}^{(N)}(\vec{r};\vec{c};\vec{x})\prod_{i=1}^{d}\widehat{\pi}_{s_{i+1},s_{i}}^{(N),a_{i+1},a_{i}}(r_{i+1},r_{i})\cdot\widehat{\pi}_{s_{d+1}}^{(N),a_{d+1}}(r_{d+1})\bigg{|}$ (4.1)
$\leq Ce^{-c(r_{1}+...+r_{d+1})}\prod_{i=1}^{d}\mathfrak{p}_{s_{i+1},s_{i}}^{a_{i+1},a_{i}}\left(r_{i+1},r_{i}\right)\cdot\mathfrak{p}_{s_{d+1}}^{a_{d+1}}(r_{d+1}).$
As in [CK24], we insert the multiplicative factor $1=\prod_{i=1}^{d+1}\left(\mathbf{1}_{r_{i}\geq 0}+\mathbf{1}_{r_{i}<0}\right)$ into the function $\mathcal{G}^{(N)}(\vec{r};\vec{c};\vec{x})$ in $\widetilde{\varphi}^{(N)}_{u,v,w,r}(\vec{c};\vec{x})$ and expand to obtain $2^{d+1}$ terms, corresponding to a choice of the atomic or continuous support at each $r_{i}$. Thus,
$\widetilde{\varphi}^{(N)}_{u,v,w,r}(\vec{c},\vec{x})$ $=\sum_{I\subset[[1,d+1]]}\sum_{r_{J}}\int_{r_{J}}\mathcal{G}(\vec{r};\vec{c};\vec{x})\prod_{i=1}^{d}\widehat{\pi}^{(N),a_{i+1},a_{i}}_{s_{i+1},s_{i}}(r_{i+1},r_{i})\cdot N^{u+v}\widehat{\pi}^{(N),a_{d+1}}_{s_{d+1}}(r_{d+1}),$
where $J$ is in the complement of $I$ in $[[1,d+1]]$ and the sum over $r_{I}$ is an $|I|$-fold sum over $r_{i}\in\widehat{S}^{(N),d}_{s_{i}}$ for $i\in I$, and the integral over $r_{J}$ is a $|J|$-fold integral of the $r$ variables with indices in $J$ over $\widehat{S}^{(N),c}$. We must show that for any choice of $I$, the corresponding sum over $r_{I}$ and integral over $r_{J}$ converge. We apply Lemmas 3.1, 3.2, 3.3 and 3.6 and the pointwise convergence of $\mathcal{G}^{(N)}(\vec{r}^{(N)};\vec{c};\vec{x})$ in Lemma 4.2 to obtain pointwise convergence of each term. To obtain dominated convergence, we need to take the dominating function from (4.1) and show that, for all $I\subset[[1,d+1]]$,
@@
which is accomplished by casework on the structure of the set $I$, see [CK24, Section 8.4].
5. Tightness and Proof of 1.1
In this section, we discuss the tightness of the laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$, and then give a full proof of 1.1. A standard result in the study of stochastic processes (see, for instance, Karatzas and Shreve [KS91]) tells us that if, for all $0\leq x,y\leq 1$ there exist $\alpha,\beta,\nu,C\in\mathbb{R}_{>0}$ such that
$\sup_{N\in\mathbb{N}}\mathbb{E}\left[|h_{u,v,w,r}^{(N)}(x)-h_{u,v,w,r}^{(N)}(y)|^{\alpha}\right]\leq C|x-y|^{1+\beta},$ (5.1)
and such that the process, evaluated at 00, has bounded supremum (which is immediate for our process since $h^{(N)}_{u,v,w,r}(0)-h^{(N)}_{u,v,w,r}(0)=0$), then the collection of laws of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ is tight on $(D[0,1],\mathcal{B}(D[0,1]))$. Since $h_{u,v,w,r}^{(N)}(0)-h_{u,v,w,r}^{(N)}(0)=0$ for all $N\in\mathbb{N}$, the proof of tightness reduces to demonstrating the second bound in (5.1).